OpenStudy (anonymous):

f(x)=x^2- x- 2 What is the axis symmetry? The Ordered pair for vertex is ?

OpenStudy (amistre64):

the axis o' symmetry is: (-) middle# ------------- 2(first#)

OpenStudy (amistre64):

-(-1) ----- = 1/2 2(1)

OpenStudy (anonymous):

? that makes no sense the axis is a single number and vertex is set of two ordered numbers? Im confused

OpenStudy (amistre64):

the axis is; x=1/2

OpenStudy (amistre64):

the vertex coords are what ever the function equals when x=1/2

OpenStudy (anonymous):

so u mean .5?

OpenStudy (anonymous):

or .05

OpenStudy (amistre64):

1 -- = .5 2

OpenStudy (anonymous):

ok ok then I just input that into the x variable to calculate the vertex order?

OpenStudy (amistre64):

yes; your vertex sits at (.5,f(.5)) so whatever the equation equals at x=.5 is gonna be your y value for your vertex coords

OpenStudy (amistre64):

the axis of symmetry is x = .5 right?

OpenStudy (anonymous):

yes ok i get it now

OpenStudy (amistre64):

i knew you would ;)

OpenStudy (anonymous):

:) Thank You